caira12:

which expression expression is equivalent to 3/4 (5z+ 16). 15/4z+16 15/4z+12 3/4z+16 3/4z+12

Champion:

Hello @caira12 For this problem, we should use something called the "distributive property". It looks something like this: 

Champion:

Here's a simple example of how the "distributive property" is true: 2*5 = 10 right? 5 also = 2+3 so 2*(2+3) = (2*2) + (2*3) = 4 + 6 = 10

Champion:

so, what we do for your problem is: (3/4)(5z + 16) = (3/4)*(5z) + (3/4)*(16) then multiply see what you get, it should end up matching one of the answers but if you are still stuck, feel free to come back and tag me with @Champion for more help :-) Good Luck!